OpenStudy (anonymous):

the domains of f(x)=5x+ 8 and g(x) 7-9x

OpenStudy (anonymous):

hw is asking for domains of f, g, fg, ff, f/g g/f

OpenStudy (anonymous):

also for (fg)(x), gf(x), f/g(x), g/f(x)

OpenStudy (anonymous):

in interval notation

OpenStudy (triciaal):

so rephrased, the question is asking what values of x will satisfy the following situations?

OpenStudy (triciaal):

f(x) all values of x will give a value for f(x) so {x: -infinity <x< infinity} f(x)=5x+ 8 and g(x) 7-9x g f(x)= 7 - 9(5 x + 8) = 7 -45 x -72 = -45x - 65 {x: x>-13/9}